please help me to solve this question
solve for 'x':
3ln2x-4=2ln2x

3In2x-4=2In2x
3In2x-2In2x=4
In2x=4
x=4/In2

Answer by ijeoma(3)

please help me to solve this question
solve for 'x':
3ln2x-4=2ln2x

3In2x-4=2In2x
3In2x-2In2x=4
In2x=4
x=4/In2